OpenStudy (anonymous):

what is the simplified form of (5)/(3y+2)-(4)/(9y^2-4)? Any help with this problem would be greatly appreciated!!!

OpenStudy (anonymous):

start with \[\frac{5}{3y+2}-\frac{4}{(3y+2)(3y-2)}\] then see that the last common multiple of the denominators is \((3y+2)(3y-2)\)

OpenStudy (anonymous):

go to \[\frac{5(3y-2)-4}{(3y+2)(3y-2)}\] and finish by multiplying out in the numerator, and combine like terms

OpenStudy (anonymous):

okay that makes sense thankyou